OpenStudy (anonymous):

what are The number of zeros of a parabola? is it always 0, 1, or 2.? just wondering?

Parth (parthkohli):

Do you know the equation of a parabola?

OpenStudy (anonymous):

No what is it parth?

OpenStudy (anonymous):

hmm..

Parth (parthkohli):

Oops

OpenStudy (anonymous):

WHY OOPS?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow y = a(x - h)^2 + b }\) I did the equation wrong. Sorry.

OpenStudy (anonymous):

its ok but im confused on what knowing the equation has to do with it?

Parth (parthkohli):

This is the main idea.

OpenStudy (anonymous):

HMMM

OpenStudy (anonymous):

^im using hmm alot now cause you say it alot lol

OpenStudy (anonymous):

BUT YEA, this is confusing..

Parth (parthkohli):

lol

Parth (parthkohli):

You just see what value can end in zero. Give me an equation. Hmm...

OpenStudy (anonymous):

lol hmm..

OpenStudy (anonymous):

y = (x - h)2 + k ?

Parth (parthkohli):

Hmm....

Parth (parthkohli):

This is a standard equation. Give me an equation.

OpenStudy (anonymous):

@dpaInc can u give him an equation ahh! LOL

OpenStudy (anonymous):

yes... because this is a quadratic (degree 2) it has at most 2 zeros.

Parth (parthkohli):

Exactly lol I was about to lead him to that statement ;) you took it away

OpenStudy (anonymous):

@ParthKohli im a girl -_- lol

Parth (parthkohli):

Really? omg

OpenStudy (anonymous):

sorry parth....:(

OpenStudy (anonymous):

LMAO yea

Parth (parthkohli):

dw.... what's up dpainc?

Parth (parthkohli):

But listen, it may even have 1 zero if it's a perfect square.

OpenStudy (anonymous):

didn't mean to cut in like that... @sktjell you can take back that medal and give it to parth...

OpenStudy (anonymous):

parth..... *dpalnc

Parth (parthkohli):

No dpainc.

Parth (parthkohli):

It's actually an i @ZhangYan

Parth (parthkohli):

It's not an L but an i @dpaInc right?

OpenStudy (anonymous):

I GAVE IT TO PARTH DPAINC :)

Parth (parthkohli):

Give it to dpainc please..

OpenStudy (anonymous):

oh ok... :)

OpenStudy (anonymous):

i gave it to him ahh! idk what to do! LOL </3

OpenStudy (anonymous):

you look like a mean owl parth.... i don't wanna mess with you!

Parth (parthkohli):

Lol let me change it

OpenStudy (anonymous):

@dpaInc how you spell your username or just this dpalnc or dpainc??

OpenStudy (anonymous):

@ParthKohli i cant believe u thought i was a guy...

OpenStudy (anonymous):

i should change my pic

OpenStudy (anonymous):

parth is correct... it is an I..

OpenStudy (anonymous):

oh ok just wondering

OpenStudy (anonymous):

haha.. it still looks like an L.. it's an "i"

OpenStudy (anonymous):

do you guys get this question?

OpenStudy (anonymous):

the equation is that of y= x^2- 4.... i'm sure you can factor that.. here's a hint: x^2-4 = x^2 - 2^2

OpenStudy (anonymous):

it asks for the factorization.

OpenStudy (anonymous):

ok so factor that equation?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):